### v1

# GraphPilot: Grounded Scene Graph Conditioning for Language-Based Autonomous Driving

Abstract
Vision-language models have recently emerged as promising planners for autonomous driving, where success hinges on topology-aware reasoning over spatial structure and dynamic interactions from multimodal input. However, existing models are typically trained without supervision that explicitly encodes these relational dependencies, limiting their ability to infer how agents and other traffic entities influence one another from raw sensor data. In this work, we bridge this gap with a novel model-agnostic method that conditions language-based driving models on structured relational context in the form of traffic scene graphs. We serialize scene graphs at various abstraction levels and formats, and incorporate them into the models via structured prompt templates, enabling a systematic analysis of when and how relational supervision is most beneficial. Extensive evaluations on the public LangAuto benchmark show that scene graph conditioning of state-of-the-art approaches yields large and persistent improvement in driving performance. Notably, we observe up to a 15.6% increase in driving score for LMDrive and 17.5% for BEVDriver, indicating that models can better internalize and ground relational priors through scene graph-conditioned training, even without requiring scene graph input at test-time. Code, fine-tuned models, and our scene graph dataset are publicly available at https://github.com/iis-esslingen/GraphPilot.
1 Introduction
Figure 1: Explicit relational grounding through scene graph conditioning. We visualize four setups, where two binary digits indicate scene graph usage during training and testing: SG00-PT (baseline pretrained, no scene graphs), SG00-FT (baseline fine-tuned without scene graphs), SG10 (scene graphs only during training), and SG11 (scene graphs during training and testing). Models trained under SG10 perform largely on par with SG11, suggesting that they internalize relational structure during training. Performance indicates (mean) driving scores, cf. Tab. 1 and Tab. 4.
Understanding complex traffic scenes remains a central challenge in autonomous driving, especially when decision-making is conditioned on natural language instructions [cui2024survey]. Modern language-based driving agents operate at the intersection of perception, language, and planning, requiring not only accurate scene understanding but also the ability to reason about spatial structure, traffic rules, and interactions among dynamic actors [xie2025vlms-ready]. While recent works have shown that vision-language models (VLMs) can map image and language inputs to plausible trajectories [shao2024lmdrive, winter2025bevdriver, renz2025simlingo, zhang2025vldrive, zhang2025adadrive], their performance in complex environments remains limited. A key limitation is that current models typically rely on representations that do not explicitly encode relational structure, forcing them to infer safety-critical interactions implicitly from dense features [zhang2025safeauto].
Scene graphs provide a structured solution to this limitation [johnson2015image, kassab2024bare]. By explicitly representing a traffic scene as a structured graph of entities and their relations, they encode not only what is present in the environment, but also how entities are related and may influence one another. In scene graphs, nodes represent objects such as vehicles, pedestrians, traffic lights, and roadway elements, while edges express spatial, semantic, and regulatory relations, e.g., which vehicle is in which lane, which traffic light controls which direction, and how nearby actors might interact. This structured relational view enhances semantic understanding and enables more informed planning [chen2025static].
Despite their growing use in indoor navigation and instruction-following tasks [gu2024conceptgraphs, rana2023sayplan, yin2024sgnav, honerkamp2024language, werby2024hierarchical, mohammadi2025more], scene graphs have only recently gained attention in outdoor autonomous driving, mainly for tasks such as risk estimation [malawade2022sg2vec, liu2023learning], and trajectory prediction [jia2023hdgt, zhang2025graphad]. However, their integration into language-based planning remains unexplored. Instead, most existing systems represent the scene using dense geometric or visual features, including image tokens [renz2025simlingo, zhang2025vldrive, zhang2025adadrive] or BEV grids [shao2024lmdrive, winter2025bevdriver, hegde2025dima], followed by generic attention mechanisms. As a result, critical relationships between entities, such as which traffic light controls a given lane or how nearby vehicles might interact, must be inferred implicitly. This gap motivates our central question: Can explicit relational structure of a traffic scene help language-based driving models internalize and reason over spatial and causal dependencies in traffic scenes?
To address this question, we introduce a model-agnostic approach for injecting structured semantic context into language-based autonomous driving systems. We construct traffic scene graphs at each time step and serialize them into human-readable formats (text, JSON, YAML), which are then provided to the language model, alongside the natural language navigation instruction and the sensory input, as shown in Fig. 1. We make this a deliberate design choice to avoid changes to the model architecture, additional training objectives, or specialized graph encoders, thereby maintaining model-agnostic applicability and ease of integration. To the best of our knowledge, this is the first work to inject a traffic scene graph as explicit context for planning in a language-based driving model.
Our experiments reveal several key insights. First, employing scene graph prompts at test-time alone improves performance over the original models. Second, training with scene graph supervision yields significant gains in driving performance, and importantly, these improvements persist even when scene graphs are omitted at test-time. This indicates that language-based planners can internalize structured relational knowledge during training and leverage it without explicit relational input at inference. Finally, we find that leaner abstractions, such as actor-only graphs with pairwise relational links, offer a strong trade-off between semantic fidelity and prompt efficiency, achieving high performance with fewer tokens.
Our main contribution is a model-agnostic approach that grounds language-based autonomous driving models in explicit relational structure through serialized traffic scene graphs, enhancing structured reasoning over spatial, regulatory, and inter-actor dependencies. We systematically evaluate our method across graph abstraction levels, serialization formats, and prompt templates with multiple state-of-the-art language-based driving models to analyze how explicit relational context influences planning behavior and performance. This reveals a novel and practical paradigm where models benefit from scene graph supervision during training but operate effectively without it at test-time, avoiding the cost and complexity of test-time scene graphs.
2 Related Work
We categorize related work into two areas: (1) language-based autonomous driving, where VLMs are used to plan from multimodal inputs and instructions, and (2) traffic scene graphs, which provide structured representations of the driving environment.
Language-Based Autonomous Driving: We review language-driven approaches for trajectory planning and decision making in autonomous driving, covering both open-loop waypoint prediction on nuScenes [caesar2020nuscenes] and closed-loop evaluation using CARLA [dosovitskiy2017carla] or NavSim [dauner2024navsim]. Several works deploy VLMs as planners, mapping sensor and a navigation command directly to a sequence of future waypoints [shao2024lmdrive, winter2025bevdriver, renz2025simlingo, zhang2025vldrive, xu2025drivegpt4v2, xie2025s4driver]. LMDrive [shao2024lmdrive] presents a VLM-based end-to-end stack that fuses camera and LiDAR inputs with a natural-language navigation command. BEVDriver [winter2025bevdriver] extends this design by strengthening the vision encoder to produce richer BEV features, improving semantic and spatial grounding. Beyond direct waypoint prediction, some approaches aim to increase reasoning fidelity by incorporating chain-of-thought within a perception-prediction-planning paradigm, accepting additional latency in exchange for interpretability and more transparent decision making [kong2025vlrdriver, zeng2025futuresightdrive, hwang2025emma, tian2024drivevlm]. Efficiency-oriented systems activate the language model only when it is expected to contribute: AdaDrive [zhang2025adadrive] learns an adaptive slow-fast collaboration that decides when and to what extent the VLM should be involved.
Other approaches decouple high-level reasoning from trajectory generation. Here, the VLM proposes subgoals, constraints, or safety checks, and a non-VLM planner converts these signals into concrete trajectories [fu2025orion, chen2025solve, han2025dme, pan2024vlp]. ORION [fu2025orion], for example, predicts planning tokens with an LLM and conditions a generative planner on these tokens to produce multi-modal trajectories. Finally, several approaches transfer the world knowledge and reasoning capabilities of VLMs into training-time supervision, removing the need for an online VLM at inference while retaining its benefits [hegde2025dima, zhang2024feedback]. In parallel, retrieval-augmented (RAG) methods condition decisions on similar scenarios or past cases to improve robustness on rare or long-tail events [yuan2024ragdriver, mei2024leapad, wen2024dilu, wang2025rad, zhang2025safeauto]. Agentic formulations expose tool APIs via function calls so the model can adaptively orchestrate perception, mapping, and planning utilities [mao2024agent].
Figure 2: Scene graph construction. Each traffic scene is represented as a structured, labeled graph capturing entities (actors, objects, structure) and their relations. We define three levels of abstraction: Full (all node types and relations), Road-Level (collapsed structural detail), and Actor-Only (actors and their pairwise interactions), enabling analysis of the trade-off between relational fidelity and prompt efficiency. Dashed nodes and edges indicate optional elements that are not always present in a given scene.
Scene Graphs in Autonomous Driving: Traffic scene graphs represent road users, roadway structure, and their relations as a graph. Research spans multiple directions: construction from sensory and map data, safety and risk estimation, semantic scene retrieval, and trajectory prediction.
On the construction side, recent work builds scene graphs from video, multi-camera perception, and HD maps, progressing from rule-based extractors to learned relation predictors and lane-topology transformers, and further to multi-agent 3D urban scene graphs [malawade2022roadscene2vec, wang2024rs2g, tian2023rsg, lv2025t2sg, tian2025query, greve2024curbsg, steinke2025curbosg, zhang2025parking]. This shift reflects a move away from hand-crafted priors toward data-driven structure that adapts to diverse traffic layouts and interaction patterns. For safety reasoning, scene graphs enable spatio-temporal modeling of interactive dynamics with multi-relational GNNs and sequence models, supporting early collision prediction and pedestrian risk assessment [malawade2022roadscene2vec, malawade2022sg2vec, liu2023learning, zhou2024hktsg]. For retrieval and downstream decision support, methods based on subgraph matching, together with hybrids that use VLM-generated descriptions, identify semantically similar scenes, mine failure patterns, and return mitigation knowledge to the planner through RAG-style augmentation [tian2023rsgsearch, tian2024rsgsearchplus, tian2025query, liao2025workzones]. Graph-based trajectory prediction approaches model interactions among agents and their surrounding map context, using graph- or transformer-based encoders to capture spatial and relational dependencies for subsequent motion forecasting [mo2022multi, zhou2022hivt, jia2023hdgt, zhang2025graphad]. Distinct from prior work, we are the first to serialize the scene graph and incorporate it directly into a language-based driving model as auxiliary context for decision making and planning.
3 Method
Our method conditions a language-based model on serialized traffic scene graphs to support structured reasoning for downstream driving decisions.
3.1 Scene Graph Construction
We first construct a scene graph for each time step by detecting nodes and extracting relations, building on [malawade2022roadscene2vec].
Formalization. We represent the scene at time $t$ as a directed, labeled multigraph $G_{t}=(V_{t},E_{t},\mathcal{R})$, where $V_{t}$ is the set of nodes, $E_{t}\subseteq V_{t}\times\mathcal{R}\times V_{t}$ the set of labeled edges, and $\mathcal{R}$ the relation vocabulary, cf. Fig. 2.
Nodes carry a unique identifier as well as a semantic class and are organized into three disjoint groups: (i) structural nodes $\mathcal{S}_{t}$, (ii) actor nodes $\mathcal{A}_{t}$, and (iii) traffic-related objects $\mathcal{O}_{t}$, so that $V_{t}=\mathcal{S}_{t}\cup\mathcal{A}_{t}\cup\mathcal{O}_{t}$. Structural nodes encode the static roadway scaffold, i.e., lanes, roads, and junctions, which provides the spatial context and connectivity for motion. Actor nodes represent road users, including the ego vehicle, which is always present, and other participants such as light vehicles (car, van, taxi, electric vehicle), heavy vehicles (truck, bus), motorcycle, bicycle, emergency vehicles, and pedestrian. These entities and their interactions define the scene dynamics that govern the ego vehicle’s driving decisions. Traffic-related objects capture regulatory context, for example traffic light, speed limit, and stop sign, which constrain admissible behavior and help anticipate required maneuvers such as stopping or yielding.
The relation vocabulary $\mathcal{R}$ is partitioned into six semantic groups. Proximity relations (safety hazard, near collision, super near, very near, near, visible) quantify distance- and risk-based interactions that are key for collision avoidance. Directional relations (direct front, side front, direct rear, side rear) describe longitudinal ordering and approach geometry that affect yielding, following, or braking decisions. Lateral relations (left of, right of) encode side-by-side context that is essential for lane keeping and safe lane changes. The hierarchical relation (is in) anchors actors and objects to their structural context (e.g., a lane “is in” a road), which localizes rules and right-of-way. Topological relations (opposes, travels to, lane change) capture connectivity and permitted motions within the lane graph. Finally, the regulatory relation (controls traffic of) ties traffic devices to the structural elements they govern.
We impose lightweight typing constraints on which node classes may be connected by which relations to ensure a compact and consistent graph. The hierarchical assignment builds the scene hierarchy by linking actors to lanes, lanes to roads, and, when present, roads to junctions via is in. The ego connects directly to its lane via is in, and ego-actor interactions are restricted to behaviorally relevant geometry, so only proximity, directional, and lateral relations are instantiated between actor pairs. Regulatory edges apply only to traffic lights and connect a traffic light node to the lanes it governs via controls traffic of. We do not add regulatory edges for speed limits or stop signs. Lane-to-lane relations capture the network structure and include topological relations (travels to, lane change, opposes) as well as lateral adjacency (to left of, to right of).
[…]
3.2 Scene Graph Serialization
Figure 3: Serialization formats. We serialize scene graphs as Text, JSON, or YAML, each encoding subject-predicate-object triplets. Text uses compact natural-language packing for brevity, JSON provides a structured and parser-friendly representation, and YAML achieves lower token counts through minimal syntax.
We serialize each graph as subject-predicate-object statements using a fixed, human-readable relation vocabulary. Fig. 3 illustrates the three formats we use, i.e., Text, JSON, and YAML, and a concrete example for each. To avoid redundancy, multiple predicates between the same ordered pair $(s,o)$ are emitted once as a multi-label statement rather than as duplicated links. When linearized, we list statements in hierarchical order with roads to junctions, lanes to roads, lane to lane connectivity, objects to lanes, and actors to lanes, and finally interactions between actors.
Text. The Text form is a compact, linearized sequence of statements separated by a vertical bar (|). It applies two packing rules: First, membership grouping aggregates multiple sources that share the same hierarchical target, yielding $s_{1},s_{2}$ is in oo”. Second, predicate merging collects all predicates that hold for a fixed ordered pair, yielding “$s$ $p_{1},p_{2},\ldots$ oo”. Statements are emitted in a fixed order with hierarchical assignments first and actor-actor interactions following.
JSON. The JSON form mirrors the graph with two arrays: nodes and links. Each node stores a stable identifier and semantic class (id, base_class). Each link stores source, target, and a labels array containing all predicates for that ordered pair. JSON is therefore packed at the pair level with one link per $(s,o)$ with possibly multiple labels, so it does not apply membership grouping. The result is a uniform, parser-friendly schema.
YAML. The YAML form is an indentation-based analogue of the JSON schema with the same two top-level collections (nodes, links) and the same fields. Like JSON, it represents one ordered pair per link and aggregates predicates in a labels list. Owing to indentation rather than braces and fewer quotation marks, YAML typically yields fewer tokens than JSON while preserving identical semantics.
3.3 Scene Graph Injection
Figure 4: Prompt templates. Three prompt templates combine scene graphs with navigation commands: V1 uses direct concatenation, V2 adds ego-role framing and section headers, and V3 introduces a structured preamble with markdown-style fencing for consistent formatting.
We instantiate three prompt templates that differ in layout and instruction framing while keeping the underlying content identical. Fig. 4 shows all three templates side by side for direct comparison.
The first template (V1) concatenates the navigation command and the scene graph in a single line without section headers. The command appears first, followed by a short cue and the serialized graph. This variant minimizes prompt overhead and matches instruction plus context compositions seen in prior work. The second template (V2) introduces explicit section headers for the scene graph and for the navigation command and begins with a reminder that the model acts as the ego vehicle. The graph is shown as a headed block and the command as a separate headed block, which makes both inputs distinct and separable when contexts become longer. The third template (V3) keeps the sectioning and wraps the scene graph in a fenced block with a format prefix such as text, json, or yaml. It also states a short Primary Objective that emphasizes adherence to the external navigation command. The fence treats the graph as read only and aims to stabilize tokenization for larger graphs or mixed formatting, and the format prefix encourages consistent parsing across representations.
3.4 Scene Graph Conditioning
We condition the model on structured relational information by including a serialized scene graph alongside the natural-language navigation command in the input prompt. In this work, we use the term conditioning to refer specifically to prompt-level supervision: the model receives explicit relational context as part of its input, but its architecture, training objective, and loss functions remain unchanged. This prompt-based conditioning allows the model to ground decision making and planning in spatial topology, regulatory context, and inter-actor interactions that would otherwise need to be inferred implicitly from visual features alone. Since scene graph-conditioning operates purely at the input level, it is model-agnostic and does not require modifying perception modules, introducing graph encoders, or defining auxiliary training losses.
4 Experiments
We evaluate how scene graph-conditioning affects the driving performance of VLM-based autonomous driving systems across multiple configurations and training setups.
4.1 Benchmark
We evaluate on the public LangAuto benchmark [shao2024lmdrive], which adapts the CARLA Leaderboard for language-based driving. The benchmark covers eight towns with diverse traffic layouts, weather, and illumination, and replaces discrete high-level commands with natural-language navigation instructions. Routes are partitioned by length into Tiny (<150m<150m), Short (150500m150-500m), and Long (>500m>500m) tracks.
Metrics. We follow the official CARLA Leaderboard metrics. Route completion (RC) measures the fraction of route distance achieved. Infraction score (IS) applies multiplicative penalties for collisions, traffic-rule violations, excessive route deviation, and timeouts, starting from 1.0 and decaying with event type and count. The driving score (DS) for a route is the product of RC and IS. We report the mean across routes as the overall benchmark score. Following the LangAuto protocol, each configuration is executed three times and results are averaged.
4.2 Baselines
We use LMDrive [shao2024lmdrive] and BEVDriver [winter2025bevdriver] as baselines. Both systems represent state-of-the-art VLM-based autonomous driving approaches and are designed to process both visual and natural language inputs. In our experiments, we use the officially released model variants: LMDrive with LLaVA-v1.5 [liu2024llava] and BEVDriver with LLaMA-7B [touvron2023llama].
Dataset Collection. The publicly available LMDrive dataset does not include scene graphs. Therefore, we extend the official LMDrive data generation pipeline with our scene graph generation, abstraction, and serialization modules (Sec. 3) so that ground-truth scene graphs are captured alongside the original modalities. The resulting data preserves the original distribution over selected routes, towns, weather, and illumination, and remains fully compatible with both LMDrive and BEVDriver.
4.3 System Configurations
For our approach, we define four scene graph usage settings, i.e., SG00, SG01, SG10, SG11, to evaluate how relational supervision affects language-based driving.
SG00 uses no scene graphs during training or test-time and serves as our main baseline. We distinguish between two variants: SG00-PT uses the original pretrained checkpoints provided by the model authors, while SG00-FT further fine-tunes these checkpoints on our collected dataset. This distinction ensures a fair comparison, as directly comparing our scene graph-conditioned models to SG00-PT would conflate the effects of additional training data with those of scene graph supervision. SG01 includes scene graphs only at test-time. This setting probes whether pretrained models can benefit from relational structure without any prior exposure to scene graphs during training. SG10 includes scene graphs during training but omits them at test-time. This measures whether models can internalize relational priors and generalize without needing explicit scene graph input during inference. SG11 includes scene graphs both during training and test-time and represents the full supervision scenario.
4.4 Implementation Details
We initialize all models from the officially released checkpoints. The language backbone is fine-tuned using LoRA adapters, while the vision encoder remains frozen. Preliminary experiments revealed that jointly optimizing both backbones or including auxiliary notice instructions during training as used in some baseline variants led to reduced closed-loop driving performance. Therefore, we omit notice instructions entirely and keep the vision encoder fixed throughout. Accordingly, we follow the default optimization settings of each baseline and do not introduce additional training losses or hyperparameter tuning strategies. All components outside the prompt remain unchanged.
4.5 Training
All models are trained on eight NVIDIA L40s GPUs. We use the same train/validation split as in the original works and train for five epochs with a global batch size of eight.
When a scene graph is used during training (SG11 or SG10), it is integrated based on the selected abstraction level, serialization format, and injection template. These choices govern the form and placement of the graph in the prompt but do not alter the model architecture or optimization procedure. The total training time per run depends primarily on the abstraction level and serialization format, ranging from 7.5 hours (Actor-Only, Text) to 13 hours (Full, JSON). Among the formats, Text is the most efficient, followed by YAML, while JSON incurs the highest overhead due to its more verbose structure. Injection template has negligible effect on training time.
4.6 Results
Table 1: Baseline performance. No scene graphs are used during training or test-time. SG00-PT refers to the official pretrained checkpoints, while SG00-FT applies further fine-tuning on our collected dataset. LMDrive degrades under SG00 fine-tuning, whereas BEVDriver improves, highlighting model-specific sensitivity to fine-tuning without relational supervision. Best model-specific results are in bold.
Method DS $\uparrow$ RC $\uparrow$ IS $\uparrow$
LMDrive (SG00-PT) 42.95 54.81 0.79
LMDrive (SG00-FT) 41.55 48.22 0.88
BEVDriver (SG00-PT) 44.70 49.70 0.90
BEVDriver (SG00-FT) 48.94 56.37 0.86
Baselines. We first establish baselines by using the official implementations and pretrained checkpoints of LMDrive and BEVDriver, referred to as SG00-PT. These results serve as primary reference points to ensure fair, environment-matched comparisons. [^1] We report reproduced results rather than official numbers, as the authors of LMDrive acknowledge that performance varies significantly depending on GPU model as reported by other authors as well (see issues 37, 52, and 56 in LMDrive repository; links omitted for review according to author guidelines). Same behaviour was observed for BEVDriver. Additionally, to isolate the effect of scene graph supervision itself, we fine-tune both models on our dataset without scene graph input at training or test-time (SG00-FT). Table 1 shows that SG00 fine-tuning shifts performance in a model-specific way: LMDrive gains in safety (higher IS) but sees a drop in RC, while BEVDriver substantially improves RC and resulting in a higher DS. These outcomes provide a neutral reference for assessing the added value of scene graph conditioning.
Table 2: Test-time-only scene graph injection (SG01). LMDrive performance using different combinations of abstraction levels, serialization formats, and injection templates. Results averaged across LangAuto Tiny, Short, and Long. Actor-Only scene graphs paired with the V3 injection template consistently perform best across formats, demonstrating that concise but structured relational input yields driving improvements even without additional training. Best mean results are highlighted in bold.
Configuration V1 V2 V3 Mean
DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$
Text Full 38.9 46.2 0.84 41.3 49.7 0.83 41.8 49.4 0.83 40.7 48.4 0.83
Road-Level 40.6 47.0 0.86 40.5 50.1 0.81 41.5 49.4 0.83 40.9 48.8 0.83
Actor-Only 40.0 48.7 0.85 43.1 51.2 0.84 44.7 53.0 0.84 42.6 51.0 0.84
JSON Full 34.6 42.2 0.84 41.5 48.7 0.84 44.6 54.6 0.83 40.2 48.5 0.84
Road-Level 34.4 43.4 0.83 41.8 51.2 0.81 46.6 56.3 0.82 41.0 50.3 0.82
Actor-Only 40.3 47.3 0.85 42.6 52.3 0.82 44.7 55.6 0.82 42.6 51.7 0.83
YAML Full 39.4 45.5 0.85 41.8 50.4 0.82 41.8 51.9 0.80 41.0 49.3 0.82
Road-Level 39.2 46.6 0.85 41.8 51.5 0.82 42.1 50.6 0.83 41.1 49.6 0.83
Actor-Only 42.8 52.0 0.84 43.6 53.2 0.82 44.7 52.7 0.84 43.7 52.6 0.83
Mean Full 37.6 44.6 0.84 41.5 49.6 0.83 42.7 52.0 0.82 40.6 48.3 0.83
Road-Level 38.1 45.7 0.84 41.4 50.9 0.81 43.4 52.1 0.83 41.0 49.6 0.83
Actor-Only 41.1 49.4 0.85 43.1 52.2 0.82 44.7 53.8 0.83 43.0 51.8 0.84
Table 3: Token statistics for scene graph variants. Mean and maximum token counts for each abstraction level and serialization format in the training set, measured using the LLaVA-v1.5 tokenizer, showing that abstraction level and serialization format substantially affect token count.
Abstraction Text JSON YAML
Mean Max Mean Max Mean Max
Full 437 1605 2905 10020 1712 5950
Road-Level 370 1576 2172 8753 1246 5052
Actor-Only 69 607 409 2591 242 1606
Test-Time-Only Scene Graph Injection (SG01). We begin by evaluating which scene graph integration strategies are most effective when applied at test-time only without additional training (SG01), to assess whether a pretrained model can benefit from relational context without having been explicitly conditioned on it during training.
Specifically, we evaluate all combinations of abstraction level, serialization format, and injection template. Table 2 reports DS, RC, and IS for LMDrive averaged over the LangAuto Tiny, Short, and Long tracks. We give mean rows and columns to support comparisons across serialization formats and injection templates, respectively. Notably, multiple test-time scene graph configurations already surpass the reproduced baseline without scene graphs (Table 1), underscoring the potential of this approach even without additional training. Two consistent trends emerge: First, injection template V3 outperforms both V2 and V1, with V1 trailing significantly. Second, Actor-Only abstractions yield the best results across all formats, indicating that minimal yet structured relational information is most beneficial when applied to pretrained models.
These results highlight the strength of lean representations and well-designed prompts, even without additional training. Actor-Only graphs in particular strike a favorable balance between semantic relevance and token budget, introducing the fewest additional tokens across all formats, as shown in Table 3. Based on these insights, we select the Actor-Only abstraction and injection templates V2 and V3 for all subsequent fine-tuning experiments, focusing on configurations that maximize prompt efficiency and clarity.
Table 4: Scene graph-conditioned fine-tuning. Comparison of SG10 and SG11 using Actor-Only graphs with V2 and V3 prompt templates. SG10 matches or exceeds SG11 in most settings, showing that models retain relational grounding and can drive effectively without test-time scene graphs. Best model-specific results are highlighted in bold.
Configuration LMDrive BEVDriver
V2 V3 V2 V3
DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$
Text Actor-Only, SG11 46.4 58.0 0.81 47.1 56.1 0.83 51.7 61.4 0.84 52.2 63.5 0.83
Actor-Only, SG10 46.8 56.2 0.84 51.8 59.4 0.86 51.7 59.8 0.87 51.5 60.0 0.86
JSON Actor-Only, SG11 43.7 55.0 0.80 48.8 55.5 0.87 53.2 62.8 0.84 43.8 51.4 0.88
Actor-Only, SG10 45.7 53.5 0.85 45.5 54.0 0.85 47.3 52.1 0.90 42.2 50.7 0.85
YAML Actor-Only, SG11 45.5 55.7 0.82 44.7 54.6 0.81 52.7 61.6 0.85 52.8 60.4 0.87
Actor-Only, SG10 47.9 57.4 0.83 46.8 54.8 0.84 56.1 64.1 0.87 48.2 55.7 0.85
Mean Actor-Only, SG11 45.2 56.2 0.81 46.9 55.4 0.84 52.5 61.9 0.85 49.6 58.4 0.86
Actor-Only, SG10 46.8 55.7 0.84 48.0 56.1 0.85 51.7 58.7 0.88 47.3 55.5 0.85
Table 5: Extended infraction analysis. Comparison of fine-tuned LMDrive models on safety and behavior sub-metrics, averaged over V3-based configurations. SG10 achieves the best overall balance, reducing collisions and deviations while avoiding the need for test-time scene graphs. Abbreviations: CP = Collision Pedestrians, CV = Collision Vehicles, CL = Collision Layout, RL = Red Light, $S$ = Stop Sign, Off = Off-Road, RD = Route Deviation, TO = Timeout, AB = Agent Blocked. Best results are highlighted in bold.
Configuration DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ CP $\downarrow$ CV $\downarrow$ CL $\downarrow$ RL $\downarrow$ $S$ $\downarrow$ Off $\downarrow$ RD $\downarrow$ TO $\downarrow$ AB $\downarrow$
SG00-PT 42.95 54.81 0.80 0.08 2.83 4.28 2.31 0.00 3.79 11.95 2.93 1.63
SG00-FT 41.55 48.22 0.88 0.00 0.85 2.93 0.10 0.00 2.86 9.94 4.31 2.51
Mean Actor-Only, SG11 47.76 56.78 0.84 0.02 1.35 2.42 1.85 0.02 3.65 5.96 3.63 3.04
Actor-Only, SG10 50.42 57.22 0.87 0.01 1.17 2.07 2.07 0.01 2.65 8.22 2.40 2.55
Scene Graph-Conditioned Fine-Tuning. Table 4 reports results for SG11 and SG10 configurations, using the best-performing abstraction (Actor-Only) and injection templates (V2 and V3). For LMDrive, SG10 consistently outperforms SG11 in DS, and V3 surpasses V2, suggesting that exposure to structured context during training enhances internal planning behavior, even when that structure is absent at test-time. The best configuration (Text, Actor-Only, V3, SG10) achieves a DS of 51.8, marking a substantial improvement over the reproduced baseline (42.95) and SG00 fine-tuning (41.55). Moreover, SG10 and SG11 significantly outperform SG01 (44.7), demonstrating that scene graph-conditioned training is more effective than test-time injection only. BEVDriver shows complementary preferences: YAML slightly outperforms Text, and V2 is favored over V3. Nevertheless, SG10 again matches or exceeds SG11, with the top configuration (YAML, Actor-Only, V2, SG10) reaching a DS of 56.1, significantly outperforming both the reproduced baseline (44.70) and SG00 (48.94). We attribute differences in optimal serialization and templates in part to the underlying language backbone (LLaVA-v1.5 in LMDrive vs. LLaMA-7B in BEVDriver), but the broader pattern holds across both: SG10 provides consistent benefits without requiring graph input at test-time.
Taken together, these findings demonstrate that scene graph supervision during training (SG10) not only improves performance but also performs comparably to those using scene graphs at both training and test-time (SG11). This highlights a central insight of our work: models can internalize structured relational information from scene graphs during training and later generalize without requiring such input at test-time. This has significant practical benefits, most notably, SG10 eliminates the need for test-time scene graph generation, which is often complex, error-prone, or costly to deploy in real-time systems. Moreover, removing scene graphs at test-time reduces prompt length, lowering computational load and memory consumption. As a result, SG10 strikes a favorable balance between leveraging structured supervision and maintaining runtime efficiency, making it especially well-suited for real-world autonomous driving applications under resource constraints.
Extended Infraction Analysis. To better understand how scene graph conditioning influences safety and behavior, we analyze detailed sub-metrics in Table 5. Rather than focusing on best-case results, we report averages over all V3-based LMDrive configurations (Text, JSON, YAML) to assess the generality of the observed effects.
Compared to the reproduced LMDrive baseline (SG00-PT), SG10 configurations yield significant reductions in critical infractions. Collisions with vehicles and layout elements drop from 2.83 and 4.28 to 1.17 and 2.07, respectively. Collisions with pedestrians are nearly eliminated. Route deviation (RD) also improves notably, dropping from 11.95 to 8.22, indicating better path adherence. These improvements suggest that scene graph injection during training leads to more cautious and structured driving, even without access to graphs at test-time. SG11 achieves similar but slightly worse results than SG10 across most sub-metrics. This could indicate a mild overreliance on explicit graph inputs during deployment or increased prompt complexity. SG00-FT improves some safety-related metrics (e.g., CV and CL), but exhibits the highest RD (9.94), implying weakened command-following precision. Overall, SG10 provides the best balance between route-following accuracy and safety, supporting the idea that scene graph-conditioned training strengthens internal decision making, even when graph inputs are unavailable at test-time.
5 Conclusion
We presented a model-agnostic approach that grounds language-based autonomous driving models in explicit relational structure through serialized traffic scene graphs. By systematically evaluating different abstraction levels, serialization formats, and injection templates, we identified lightweight and effective strategies that improve driving performance without requiring architectural changes or additional training objectives. Among all configurations, Actor-Only abstractions, which include only dynamic agents and their pairwise interactions, provide a strong balance between semantic relevance and token efficiency. Our experiments further revealed that scene graph prompts enhance planning behavior even when applied only at test-time to pretrained models. More importantly, training with scene graph supervision yields additional consistent performance gains that persist even when scene graphs are removed at test-time. These findings indicate that structural priors can foster more grounded and consistent reasoning in language-based autonomous driving systems.
Limitations. While our approach improves planning through explicit relational supervision, it relies on access to accurate scene graphs during training. Although we show that test-time graph input is not required, generating high-quality scene graphs for training remains a challenge, especially in unstructured or map-less environments. Moreover, we currently inject the serialized scene graph purely at the input level. While this keeps the approach model-agnostic and easy to adopt, future work may explore integrating relational structure more directly within the model backbone to further strengthen visual–relational alignment.